# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v1

@@ 6 Conclusion
The appendix is structured as follows: We first present the derivations excluded from the main paper due to space limitation in Section A. Section B introduces the concept and examples of higher-order networks. Additional explanations on related work are provided in Section C. Section D details the generation process, including the architecture of the proposed denoising network, as well as the training and sampling procedures. Computational efficiency is discussed in Section E. Section F outlines the experimental setup, and Section G concludes with visualizations of the generated samples.
@@ Appendix A Formal Statements and Proofs
This section presents the formal statements of key theoretical results and their detailed derivations. We will recall and more precisely state the propositions before presenting the proof.
@@ A.1 Diffusion Bridge Process
In the following, we derive the Generalized Ornstein-Uhlenbeck (GOU) bridge process using Doob’s $h$-transform (Doob & Doob, 1984) and analyze its relationship with the Brownian bridge process.
Recall that the generalized Ornstein-Uhlenbeck (GOU) process is the time-varying OU process. It is a stationary Gaussian-Markov process whose marginal distribution gradually tends towards a stable mean and variance over time. The GOU process $\mathbb{Q}$ is generally defined as follows (Ahmad, 1988; Luo et al., 2023b):
$\mathbb{Q}:\mathrm{d}\bm{G}_{t}=\theta_{t}\left(\bm{\mu}-\bm{G}_{t}\right)\mathrm{d}t+g_{t}\mathrm{d}\bm{W}_{t},$ (13)
where $\bm{\mu}$ is a given state vector, $\theta_{t}$ denotes a scalar drift coefficient and $g_{t}$ represents the diffusion coefficient. At the same time, we require $\theta_{t},g_{t}$ to satisfy the specified relationship $2\sigma^{2}=g_{t}^{2}/\theta_{t}$, where $\sigma^{2}$ is a given constant scalar. As a result, its transition probability possesses a closed-form analytical solution:
@@
[…]
Proposition 3.2.
Let $\bm{G}_{t}$ evolve according to the generalized OU process in Eq. 8, subject to the terminal conditional $\bm{\mu}=\bm{G}_{\tau_{k}}$. The conditional marginal distribution $p(\bm{G}_{t}\mid\bm{G}_{\tau_{k}})$ then evolves according to the following SDE:
@@
The conditional transition probability $p(\bm{G}_{t}\mid\bm{G}_{\tau_{k-1}},\bm{G}_{\tau_{k}})$ has analytical form as follows:
@@
To simplify the notion, in the $k$-th generation step, we adopt the following conventions: $T=\tau_{k}$, $\mathbf{x}_{t}=\bm{G}_{t}^{(k)}$, $0=\tau_{k-1}$, $\mathbf{x}_{0}=\bm{G}_{\tau_{k-1}}$, $\mathbf{x}_{T}=\bm{G}_{\tau_{k}}$.
@@
Then the Doob’s $h$-transform yields the representation of an endpoint $\mathbf{x}_{T}$ conditioned process defined by the following SDE:
@@
Given that the joint distribution of $[\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{x}_{T}]$ is multivariate normal, the conditional distribution $p(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})$ is also Gaussian:
@@
Note that the generalized OU bridge process, also referred to as the conditional GOU process, has been studied theoretically in previous works (Salminen, 1984; Heng et al., 2021; Yue et al., 2024). However, we are the first to demonstrate its effectiveness in explicitly learning higher-order structures within the graph generation process.
@@
Assume $\theta_{t}=\theta$ is a constant that tends to zero, we obtain
@@
Consequently, in the limit $\theta_{t}\rightarrow 0$, the generalized OU bridge process described in Eq. 15 can be modelled by the following SDE:
@@
In contrast to the generalized OU bridge process in Eq. 15, the evolution of the Brownian bridge is fully determined by the noise schedule $g_{k,t}$, resulting in a simpler SDE representation. However, this constraint in the Brownian bridge reduces the flexibility in designing the generative process.
@@
This equation describes a pure diffusion process without drift, making it a specific instance of the generalized OU process in Eq. 8.
A.2 Proof of Proposition 3.3
@@
Definition A.1 ($\beta$-smooth).
@@
Next, we show that the proposed framework has a smaller smoothness constant than the classical one-step model. Therefore, we focus exclusively on the spectral component $||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||_{2}^{2}$ from the full loss function in Eq. 11, as the feature-related part of the loss function in HOG-Diff aligns with that of the classical framework. For simplicity, we use the notation ¯$\bar{\ell}(\bm{\theta})=||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||^{2}=||\bm{s}_{\bm{\theta}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t})||^{2}$ as the feature-related part of the loss.
Next, we verify that ¯$\bar{\ell}(\bm{\theta})$ is $\beta$-smooth under the assumptions given. Notice that the gradient of the loss function is given by:
¯$\nabla\bar{\ell}(\bm{\theta})=2\mathbb{E}\left[(\bm{s}_{\bm{\theta}}(\mathbf{x})-\nabla\log p(\mathbf{x}))^{\top}\nabla_{\bm{\theta}}s_{\bm{\theta}}(\mathbf{x})\right]$ (34)
@@
¯$\begin{split}&\|\nabla\bar{\ell}(\bm{\theta}_{1})-\nabla\bar{\ell}(\bm{\theta}_{2})\|=2\left\|\mathbb{E}\left[(\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-\nabla\log p(\mathbf{x}))^{\top}\nabla\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-(\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})-\nabla\log p(\mathbf{x}))^{\top}\nabla\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})\right]\right\|\\ &\leq 2\mathbb{E}[\|\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})\|\cdot\|\nabla\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})\|+\|\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})-\nabla\log p(\mathbf{x})\|\cdot\|\nabla\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-\nabla\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})\|].\end{split}$ (35)
@@
¯$\begin{split}\|\nabla\bar{\ell}(\bm{\theta}_{1})-\nabla\bar{\ell}(\bm{\theta}_{2})\|\ &\leq 2\mathbb{E}\left[C_{1}\beta_{\bm{s}_{\bm{\theta}}}\|\bm{\theta}_{1}-\bm{\theta}_{2}\|+C_{2}\beta_{\nabla\bm{s}_{\bm{\theta}}}\|\bm{\theta}_{1}-\bm{\theta}_{2}\|\right]\\ &=2(\beta_{\bm{s}_{\bm{\theta}}}C_{1}+C_{2}\beta_{\nabla\bm{s}_{\bm{\theta}}})\|\bm{\theta}_{1}-\bm{\theta}_{2}\|.\end{split}$ (36)
To satisfy the $\beta$-smooth of ¯$\bar{\ell}(\bm{\theta})$, we require that
@@
Consider the loss function of classical generative models goes as: ¯$\bar{\ell}(\bm{\varphi})=\mathbb{E}||\bm{s}_{\bm{\varphi}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}_{t}}q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})||^{2}$. To demonstrate that the proposed framework converges faster to the target distribution compared to the classical one-step generation framework, it suffices to show that: $\beta_{\nabla p_{t|\tau_{k-1}}}\leq\beta_{\nabla q_{t|0}}$ and $\beta_{\nabla^{2}p_{t|\tau_{k-1}}}\leq\beta_{\nabla^{2}q_{t|0}}$.
@@
$\begin{split}||\nabla_{\mathbf{x}^{\prime}}^{k}p(\mathbf{x}^{\prime}_{1})-\nabla_{\mathbf{x}^{\prime}}^{k}p(\mathbf{x}^{\prime}_{2})||&=||\nabla_{\mathbf{x}^{\prime}}^{k}\int\left(q(\mathbf{x}_{1}^{\prime}-\epsilon)-q(\mathbf{x}_{2}^{\prime}-\epsilon)\right)\pi(\epsilon)\mathop{}\!\mathrm{d}{\epsilon}||\\ &\leq\int||\nabla_{\mathbf{x}^{\prime}}^{k}q(\mathbf{x}_{1}^{\prime}-\epsilon)-\nabla_{\mathbf{x}^{\prime}}^{k}q(\mathbf{x}_{2}^{\prime}-\epsilon)||\pi(\epsilon)\mathop{}\!\mathrm{d}{\epsilon}\\ &\leq||\nabla_{\mathbf{x}^{\prime}}^{k}q(\mathbf{x}^{\prime})||_{\mathrm{lip}}(\mathbf{x}_{1}^{\prime}-\mathbf{x}_{2}^{\prime})\int\pi(\epsilon)\mathop{}\!\mathrm{d}{\epsilon}\\ &\leq||\nabla_{\mathbf{x}^{\prime}}^{k}q(\mathbf{x}^{\prime})||_{\mathrm{lip}}(\mathbf{x}_{1}^{\prime}-\mathbf{x}_{2}^{\prime}).\end{split}$ (38)
Hence, $||\nabla_{\mathbf{x}^{\prime}}^{k}\log p(\mathbf{x}^{\prime})||_{\mathrm{lip}}\leq||\nabla_{\mathbf{x}^{\prime}}^{k}\log q(\mathbf{x}^{\prime})||_{\mathrm{lip}}$.
@@
where $\alpha(0)=C^{2}\ell^{(1)}(\bm{\theta})\int_{0}^{\tau_{1}}g_{1,s}^{4}\mathop{}\!\mathrm{d}{s}+C\mathcal{E}(\tau_{1})\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}$, $\gamma(s)=C^{2}g_{1,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{1,s}\|_{\mathrm{lip}}^{2}$, and $h_{1,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{1}}}-1}\right)$. Furthermore, we can derive that the reconstruction error bound of HOG-Diff is sharper than classical graph generation models.
@@
Let $h(s,\tau)=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau}}-1}\right)$, $a(\tau)=\int_{0}^{\tau}h(s,\tau)^{2}\mathop{}\!\mathrm{d}{s}$, and $b(\tau)=\int_{0}^{\tau}\|h(s,\tau)\|_{\operatorname{lip}}^{2}\mathop{}\!\mathrm{d}{s}$ Since $\tau_{1}\leq T$, it follows that $\mathcal{E}(\tau_{1})\leq\mathcal{E}^{\prime}(T)$. Additionally, by Prop. 3.3, $\ell(\cdot)$ converges exponentially in the score-matching process. Therefore, to prove $\mathcal{E}(0)\leq\mathcal{E}^{\prime}(0)$, it suffices to show that both $a(\tau)$ and $b(\tau)$ are increasing functions.
@@
Since $h(\tau,\tau)\rightarrow 0$, we can derive that $a^{\prime}(\tau)>0$ and $b^{\prime}(\tau)>0$. This implies $\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}\leq\int_{0}^{T}h_{s}^{2}\mathop{}\!\mathrm{d}{s}$ and $\int_{0}^{\tau_{1}}\gamma(s)\mathop{}\!\mathrm{d}{s}\leq\int_{0}^{T}\gamma^{\prime}(s)\mathop{}\!\mathrm{d}{s}$. Combining these inequalities, we can finally conclude $\mathcal{E}(0)\leq\mathcal{E}^{\prime}(0)$. Therefore, HOG-Diff provides a sharper reconstruction error bound than the classical graph generation framework. ∎
Appendix B Higher-order Networks
Graphs are elegant and useful abstractions for modelling irregular relationships in empirical systems, but their inherent limitation to pairwise interactions restricts their representation of group dynamics (Battiston et al., 2020; Xiao & Blanco, 2022). For example, cyclic structures like benzene rings and functional groups play a holistic role in molecular networks; densely interconnected structures, like simplices, often have a collective influence on social networks; and functional brain networks exhibit higher-order dependencies. To address this, various topological models have been employed to describe data in terms of its higher-order relations, including simplicial complexes (Huang et al., 2024), cell complexes (Bodnar et al., 2021), and combinatorial complexes (Hajij et al., 2023). As such, the study of higher-order networks has gained increasing attention for their capacity to capture higher-order interactions (Papamarkou et al., 2024; Zeng et al., 2024a).

### v2

@@ Acknowledgments
The appendix is structured as follows: We first present the derivations excluded from the main paper due to space limitations in Section A. Additional explanations of related work are provided in Section B. Section C details the generation process, including the spectral diffusion framework, the architecture of the proposed score network, and the training and sampling procedures. Computational efficiency is discussed in Section D. Section E outlines the experimental setup, and Section F reports additional experimental results, covering the impact of diffusion domain choice, scalability to large graphs, standard deviation analysis, and visualizations of the generated samples. Section G concludes with limitations.
@@ Appendix A Formal Statements and Proofs
This section presents the formal statements of key theoretical results along with their detailed derivations. We will recall and more precisely state the theoretical claims before presenting the proof.
@@ A.1 Diffusion Bridge Process
Recall that the generalized Ornstein-Uhlenbeck (GOU) process, also known as the time-varying OU process, is a stationary Gaussian-Markov process whose marginal distribution gradually tends towards a stable mean and variance over time. The GOU process is generally defined as follows (Ahmad, 1988; Luo et al., 2023b):
$\mathrm{d}\bm{G}_{t}=\theta_{t}\left(\bm{\mu}-\bm{G}_{t}\right)\mathrm{d}t+g_{t}\mathrm{d}\bm{W}_{t},$ (11)
where $\bm{\mu}$ is a given state vector, $\theta_{t}$ denotes a scalar drift coefficient, and $g_{t}$ represents the diffusion coefficient. Additionally, we assume the relation $g_{t}^{2}/\theta_{t}=2\sigma^{2}$, where $\sigma^{2}$ is a given constant scalar. As a result, its transition probability possesses a closed-form analytical solution:
@@
Here, $\bar{\theta}_{s:t}=\int_{s}^{t}\theta_{z}dz$. When $s=0$, we write $\bar{\theta}_{t}:=\bar{\theta}_{0:t}$ for notation simplicity.
The Doob’s $h$-transform can modify an SDE to pass through a specified endpoint (Doob & Doob, 1984). When applied to the GOU process, it eliminates variance in the terminal state by driving the diffusion process toward a Dirac distribution centered at $\bm{G}_{\tau_{k}}$ (Heng et al., 2021; Yue et al., 2024), making it well-suited for stochastic modelling with terminal constraints.
In the following, we derive the generalized Ornstein–Uhlenbeck (GOU) bridge process using Doob’s $h$-transform (Doob & Doob, 1984), and subsequently examine its relationship with the Brownian bridge process.
Generalized Ornstein–Uhlenbeck (GOU) bridge
Let $\bm{G}_{t}$ evolve according to the generalized OU process in Eq. 4, subject to the terminal conditional $\bm{\mu}=\bm{G}_{\tau_{k}}$. Applying Doob’s $h$-transform (Doob & Doob, 1984), we can derive the GOU bridge process as follows:
@@
The conditional transition probability $p(\bm{G}_{t}\mid\bm{G}_{\tau_{k-1}},\bm{G}_{\tau_{k}})$ admits an analytical expression:
@@
Without loss of generality, consider one generation interval $\tau_{k-1},\tau_{k}]$ and denote $T=\tau_{k}$, $\mathbf{x}_{t}=\bm{G}_{t}^{(k)}$, $0=\tau_{k-1}$, and endpoints $\mathbf{x}_{0}=\bm{G}_{\tau_{k-1}}$, $\mathbf{x}_{T}=\bm{G}_{\tau_{k}}$.
@@
Following the approach in Yue et al. (2024), applying the Doob’s $h$-transform yields the representation of an endpoint $\mathbf{x}_{T}$ conditioned process defined by the following SDE:
@@
Given that the joint distribution of $\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{x}_{T}]$ is multivariate normal, the conditional distribution $p(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})$ is also Gaussian:
@@
Note that the GOU bridge process, also referred to as the conditional GOU process, has been studied theoretically in previous works (Salminen, 1984; Heng et al., 2021; Yue et al., 2024). However, we are the first to demonstrate its effectiveness in explicitly learning higher-order structures within the graph generation process.
@@
Assuming $\theta_{t}=\theta$ is a constant that tends to zero, we obtain
@@
Consequently, in the limit $\theta_{t}\rightarrow 0$, the GOU bridge process described in Eq. 13 can be modelled by the following SDE:
@@
In contrast to the GOU bridge process in Eq. 13, the evolution of the Brownian bridge is fully determined by the noise schedule $g_{k,t}$, resulting in a simpler SDE representation. However, this constraint in the Brownian bridge reduces the flexibility in designing the generative process.
@@
This equation describes a pure diffusion process without drift, making it a specific instance of the GOU process.
A.2 Proof of Theorem 3
@@
Definition 5 ($\beta$-smooth).
@@
Next, we show that the proposed framework has a smaller smoothness constant than the classical one-step model. Therefore, we focus exclusively on the spectral component $||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||_{2}^{2}$ from the full loss function in Eq. 8, as the feature-related part of the loss function in HOG-Diff aligns with that of the classical framework. For simplicity, we use the notation $\bar{\ell}(\bm{\theta})=||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||^{2}=||\bm{s}_{\bm{\theta}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t})||^{2}$ as the feature-related part of the loss.
Next, we verify that $\bar{\ell}(\bm{\theta})$ is $\beta$-smooth under the assumptions given. Notice that the gradient of the loss function is given by:
$\nabla\bar{\ell}(\bm{\theta})=2\mathbb{E}\left[(\bm{s}_{\bm{\theta}}(\mathbf{x})-\nabla\log p(\mathbf{x}))^{\top}\nabla_{\bm{\theta}}s_{\bm{\theta}}(\mathbf{x})\right]$ (32)
@@
$\begin{split}&\|\nabla\bar{\ell}(\bm{\theta}_{1})-\nabla\bar{\ell}(\bm{\theta}_{2})\|=2\left\|\mathbb{E}\left[(\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-\nabla\log p(\mathbf{x}))^{\top}\nabla\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-(\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})-\nabla\log p(\mathbf{x}))^{\top}\nabla\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})\right]\right\|\\ &\leq 2\mathbb{E}[\|\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})\|\cdot\|\nabla\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})\|+\|\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})-\nabla\log p(\mathbf{x})\|\cdot\|\nabla\bm{s}_{\bm{\theta}_{1}}(\mathbf{x})-\nabla\bm{s}_{\bm{\theta}_{2}}(\mathbf{x})\|].\end{split}$ (33)
@@
$\begin{split}\|\nabla\bar{\ell}(\bm{\theta}_{1})-\nabla\bar{\ell}(\bm{\theta}_{2})\|\ &\leq 2\mathbb{E}\left[C_{1}\beta_{\bm{s}_{\bm{\theta}}}\|\bm{\theta}_{1}-\bm{\theta}_{2}\|+C_{2}\beta_{\nabla\bm{s}_{\bm{\theta}}}\|\bm{\theta}_{1}-\bm{\theta}_{2}\|\right]\\ &=2(\beta_{\bm{s}_{\bm{\theta}}}C_{1}+C_{2}\beta_{\nabla\bm{s}_{\bm{\theta}}})\|\bm{\theta}_{1}-\bm{\theta}_{2}\|.\end{split}$ (34)
To satisfy the $\beta$-smooth of $\bar{\ell}(\bm{\theta})$, we require that
@@
Consider the loss function of classical generative models goes as: $\bar{\ell}(\bm{\varphi})=\mathbb{E}||\bm{s}_{\bm{\varphi}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}_{t}}q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})||^{2}$. To demonstrate that the proposed framework converges faster to the target distribution compared to the classical one-step generation framework, it suffices to show that: $\beta_{\nabla p_{t|\tau_{k-1}}}\leq\beta_{\nabla q_{t|0}}$ and $\beta_{\nabla^{2}p_{t|\tau_{k-1}}}\leq\beta_{\nabla^{2}q_{t|0}}$.
@@
[…]
@@ A.3 Proof of Theorem 4
where $\alpha(0)=C^{2}\ell^{(1)}(\bm{\theta})\int_{0}^{\tau_{1}}g_{1,s}^{4}\mathop{}\!\mathrm{d}{s}+C\mathcal{E}(\tau_{1})\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}$, $\gamma(s)=C^{2}g_{1,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{1,s}\|_{\mathrm{lip}}^{2}$, and $h_{1,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{1}}}-1}\right)$. Furthermore, we can derive that the reconstruction error bound of HOG-Diff is sharper than that of classical graph generation models.
@@
Let $h(s,\tau)=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau}}-1}\right)$, $a(\tau)=\int_{0}^{\tau}h(s,\tau)^{2}\mathop{}\!\mathrm{d}{s}$, and $b(\tau)=\int_{0}^{\tau}\|h(s,\tau)\|_{\operatorname{lip}}^{2}\mathop{}\!\mathrm{d}{s}$ . Since $\tau_{1}\leq T$, it follows that $\mathcal{E}(\tau_{1})\leq\mathcal{E}^{\prime}(T)$. Additionally, by Theorem 3, $\ell(\cdot)$ converges exponentially in the score-matching process. Therefore, to prove $\mathcal{E}(0)\leq\mathcal{E}^{\prime}(0)$, it suffices to show that both $a(\tau)$ and $b(\tau)$ are increasing functions.
@@
Since $h(\tau,\tau)\rightarrow 0$, we can derive that $a^{\prime}(\tau)>0$ and $b^{\prime}(\tau)>0$. This implies $\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}\leq\int_{0}^{T}h_{s}^{2}\mathop{}\!\mathrm{d}{s}$ and $\int_{0}^{\tau_{1}}\gamma(s)\mathop{}\!\mathrm{d}{s}\leq\int_{0}^{T}\gamma^{\prime}(s)\mathop{}\!\mathrm{d}{s}$. Combining these inequalities, we can finally conclude $\mathcal{E}(0)\leq\mathcal{E}^{\prime}(0)$. Therefore, HOG-Diff provides a sharper reconstruction error bound than classical graph generation frameworks. ∎
Appendix B Additional Explanation on Related Works
B.1 Higher-order Networks
Graphs are elegant and useful abstractions for modeling irregular relationships in empirical systems, transforming unstructured data into analyzable representations. However, their inherent limitation to pairwise interactions restricts their representation of group dynamics (Battiston et al., 2020). For example, cyclic structures like benzene rings and functional groups play a holistic role in molecular networks; densely interconnected structures, like simplices, often have a collective influence on social networks; and functional brain networks exhibit higher-order dependencies. To address this, various topological models have been employed to describe data in terms of its higher-order relations, including simplicial complexes, cell complexes, and combinatorial complexes (Hajij et al., 2023). As such, the study of higher-order networks has gained increasing attention for their capacity to capture higher-order interactions, with broad applications across domains such as social network analysis (Zeng et al., 2024), graph signal processing (Roddenberry et al., 2022; Sardellitti et al., 2021), and topological deep learning (Bodnar et al., 2021; Huang et al., 2024; Papamarkou et al., 2024).